### v3

# Static Spherical Vacuum Solution to Bumblebee Gravity with Time-like VEVs

Abstract
The static spherical vacuum solution in a bumblebee gravity model where the bumblebee field $B_{\mu}$ has a time-like vacuum expectation value $b_{\mu}$ is studied. We show that in general curved space-time solutions are not allowed and only the Minkowski space-time exists. However, it is surprising that non-trivial solutions can be obtained so long as a unique condition for the vacuum expectation $b^{2}\equiv-b^{\mu}b_{\mu}=2/\kappa$, where $\kappa=8\pi G$, is satisfied. We argue that naturally these solutions are not stable since quantum corrections would invalidate the likely numerical coincidence, unless there are some unknown fine-tuning mechanisms preventing any deviation from this condition. Nevertheless, some properties of these novel but peculiar solutions are discussed, and we show that the extremal Reissner-Nordström solution is a limit of one of our solutions.
I Introduction
General Relativity (GR) and the Standard Model (SM) of particle physics are the most successful theories describing all four fundamental forces of nature. However, there are theoretical tensions between GR and SM, and to reconcile them, several candidates of quantum gravity (QG) theories have already been proposed. Generally, the onset of significant effects of QG is expected to happen at the Planck scale ($E_{Pl}\sim 10^{19}~{}\text{GeV}$), which is far beyond our reach for current experiments. Although direct detection of QG effects seems to be unlikely at present, it is suggested that there exists the possibility that certain kinds of remnant signals of QG could be observed at energy scales much lower than the Planck scale. One of such signals is the violation of Lorentz invariance.
In recent decades, increasing interest has been directed toward possible violations of Lorentz symmetry, driven by attempts to formulate a consistent theory of quantum gravity and to understand potential deviations from GR at high energies. Among various approaches, effective field theories incorporating spontaneous Lorentz violation have proven especially fruitful. In this context, Bumblebee gravity has emerged as a minimal yet nontrivial extension of GR, wherein a vector field acquires a nonzero vacuum expectation value (VEV), leading to spontaneous breaking of local Lorentz invariance and diffeomorphism invariance in a controlled manner [1, 2, 3, 4, 5].
The Bumblebee model typically introduces a vector field $B_{\mu,}$ coupled non-minimally to gravity and governed by a potential $V(B_{\mu}B^{\mu}\pm b^{2})$, which determines the vacua of the theory at the classical level. The vacua, determined by the vacuum expectation $\langle B^{\mu}\rangle$ such that $V|_{B_{\mu}=\langle B_{\mu}\rangle}$, do not transform as scalars, thus signaling the spontaneously breaking of Lorentz symmetry. This spontaneous Lorentz symmetry breaking results in modifications to the Einstein field equations, leading to potentially observable signatures in gravitational phenomena. In previous studies, an exact Schwarzschild-like solution in bumblebee gravity is proposed [6]. Also, an exact Kerr-like solution is found [7]. But for both of the solutions, the configurations of the bumblebee field only admit space-like VEVs, i.e., $\langle B^{\mu}B_{\mu}\rangle>0$. Similar results are also obtained and analyzed in detail within metric-affine bumblebee models [8, 9, 10]. The purpose of this work is to investigate the solution to the Bumblebee gravity model with a time-like vacuum expectation value of the bumblebee field.
This literature is organized as follows. In Sec. II, we briefly introduce the action and the equation of motion for bumblebee gravity. In Sec. III, we try to solve the equations of motion in a static spherical field configuration with the bumblebee field obtaining time-like VEVs, and we find that for general VEVs of this type there is no solution for the equations, unless $b=\sqrt{2/\kappa}$. In Sec. IV, we solve the equations when $b=\sqrt{2/\kappa}$, and find two kinds of non-trivial solutions. Sec. V is the discussion and the summary. In this literature, we will adopt the metric signature ($+$)(-$+$)( - + + + ) and also all quantities involved are expressed in natural units $(\hbar=c=1)$.
II Brief Introduction of Bumblebee gravity
The action of Bumblebee gravity can be expressed as [6]  [^1] We have omitted additional terms proportional to the cosmological constant $\Lambda$, $B^{\mu}B_{\mu}R$, $\nabla_{\mu}B_{\nu}\nabla^{\mu}B^{\nu}$ and $(\nabla_{\mu}B^{\mu})^{2}$
$S=$ $\int d^{4}x\sqrt{-g}\left(\frac{1}{2\kappa}R+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B^{\mu\nu}B_{\mu\nu}-V\right)$
$+S_{\mathrm{m}},$ (1)
where $g$ is the determinant of the metric $g_{\mu\nu}$, the constant $\kappa\equiv 8\pi G$ with $G$ being the gravitational constant, $S_{\mathrm{m}}$ represents the action for matter fields of no interest in this work, $B_{\mu}$ is the bumblebee field, and the field strength tensor is $B_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}$. In bumblebee theories, the potential $V$ is selected to provide a non-vanishing VEV for $B_{\mu}$, and could have the following general functional form
$V\equiv V(B^{\mu}B_{\mu}+sb^{2}),$ (2)
where $b$ is a positive real constant, and $s=\pm 1$ to determine whether the expection value of $B_{\mu}$ is time-like or space-like. In the literature, it is usually assumed that $V$ has (at least one of) its minimum/maximum at 00, thus  [^2] The condition for $V^{\prime\prime}(0)$ plays no role in this work.
$V(0)=0,\ \text{and}\ V^{\prime}(0)=0.$ (3)
The VEV of the bumblebee field is determined when $V(B^{\mu}B_{\mu}+sb^{2})=0$, implying that
$B^{\mu}B_{\mu}+sb^{2}=0,$ (4)
The above equation provides a non-null vacuum expectation value
$\langle B^{\mu}\rangle=b^{\mu},$ (5)
where $b_{\mu}b^{\mu}+sb^{2}=0$.
We are interested in the vacuum solution, namely $(T_{m})_{\mu\nu}=0$. In this work, we consider the case that $B_{\mu}$ admits VEV as $b_{\mu}$ and there is no cosmological constant, so for $V$ we have
$\left.V\right|_{B_{\mu}=b_{\mu}}$ $=0,$ (6)
$\left.\frac{dV}{d(B^{\lambda}B_{\lambda})}\right|_{B_{\mu}=b_{\mu}}$ $=0.$
Provided with the above condition for $V$ and $B_{\mu}$ replaced exactly by its VEV $b_{\mu}$, the equation of motions for $g_{\mu\nu}$ is
$R_{\mu\nu}-\frac{1}{2}R\,g_{\mu\nu}-\kappa T^{B}_{\mu\nu}=0,$ (7)
where
$T_{\mu\nu}^{B}=$ $-B_{\mu\alpha}B_{\nu}^{\alpha}-\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}g_{\mu\nu}$ (8)
$+\frac{\xi}{\kappa}\left[\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}\right]$
$+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)$
$-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha}B^{\beta}\right).$
Tracing out the equation for $g_{\mu\nu}$, we get $R=-\kappa T^{B}$, and thus the above equation can be expressed as
$R_{\mu\nu}-\kappa(T^{B}_{\mu\nu}-\frac{1}{2}g_{\mu\nu}T^{B})=0.$ (9)
For $B_{\mu}$, the equation of motion is
$\frac{\xi}{\kappa}B^{\mu}R_{\mu\nu}+\nabla^{\mu}B_{\mu\nu}=0.$ (10)
In fact, Eq. (9) and Eq. (10) are the same equations of motion for a massless vector field non-minimally coupled to gravity, i.e., action Eq. (1) with $V\equiv 0$. But in the bumblebee model, one more constraint $B_{\mu}B^{\mu}+sb^{2}=0$ is added. Luckily, the exact Schwarzschild-like solution with space-like VEV for $B_{\mu}$, explored in Ref. [6], satisfies all of the equations of motion and the constraint. In general, the existence of a solution that satisfies all of the equations and the constraint simultaneously is unlikely. And this is what we will discuss in the next section.
III Static spherical Vacuum solution for bumblebee gravity with general time-like VEVs
Here we consider the time-like case and fix $s=1$ in Eq. (2) in the following. According to the Birkhoff theorem, we adopt the following metric
$g_{\mu\nu}=\mathrm{diag}\left(-e^{2\alpha},e^{2\beta},r^{2},r^{2}\sin^{2}\theta\right),$ (11)
where $\alpha$ and $\beta$ are functions of $r$. We consider a time-like ground $b_{\mu}$ as
$b_{\mu}=(b_{t}(r),0,0,0).$ (12)
By $b_{\mu}b^{\mu}+b^{2}=0$, we have
$b_{t}(r)=be^{\alpha(r)}.$ (13)
Substitute the above equations into Eq. (9) and Eq. (10) and define $\ell=\xi b^{2}$, we have the following expressions to be zero for the gravity sector
$\rm{EQ}_{tt}=$ $+(\ell-2)r\alpha^{\prime\prime}(r)+(\kappa b^{2}-2+\ell)r\alpha^{\prime}(r)^{2}$ (14a)
$-(\ell-2)\alpha^{\prime}(r)\left(r\beta^{\prime}(r)-2\right),$
$\rm{EQ}_{rr}=$ $-(\ell+2)r\alpha^{\prime\prime}(r)+(\kappa b^{2}-2-\ell)r\alpha^{\prime}(r)^{2}$ (14b)
$+4\beta^{\prime}(r)+\alpha^{\prime}(r)\left((\ell+2)r\beta^{\prime}(r)-2\ell\right),$
$\rm{EQ}_{\theta\theta}=$ $-\ell r^{2}\alpha^{\prime\prime}(r)-(\kappa b^{2}+\ell)r^{2}\alpha^{\prime}(r)^{2}$ (14c)
$+r\alpha^{\prime}(r)(\ell r\beta^{\prime}(r)-2(1+\ell))$
$+2(r\beta^{\prime}(r)+e^{2\beta(r)}-1),$
$\rm{EQ}_{\phi\phi}=$ $\sin^{2}\theta\cdot\rm{EQ}_{\theta\theta},$ (14d)
and
$\rm{EQ}^{B}_{t}=$ $-(\kappa b^{2}-\ell)r\alpha^{\prime\prime}(r)+\ell r\alpha^{\prime}(r)^{2}$ (15)
$+(\kappa b^{2}-\ell)\alpha^{\prime}(r)(r\beta^{\prime}(r)-2)$
for $B_{\mu}$ sector. Notice that by applying
$(\kappa b^{2}-\ell)\rm{EQ}_{tt}+(\ell-2)\rm{EQ}^{B}_{t}=0,$ (16)
we have the following relation
$\kappa b^{2}(\kappa b^{2}-2)r\alpha^{\prime}(r)^{2}=0.$ (17)
So for general cases with $b\neq 0$ and $b\neq\sqrt{2/\kappa}$, we have $\alpha^{\prime}(r)=0$. Substitute it into Eq. (14b), we have $\beta^{\prime}(r)=0$. Thus, there is no non-trivial static spherical solution for bumblebee gravity with general time-like VEVs. This is one of the main results of this work. Next we focus on solutions with the condition $b=\sqrt{2/\kappa}$ satisfied.
IV Solution for bumblebee gravity with VEV $b=\sqrt{2/\kappa}$
Notice that by applying
$\frac{r}{4}\left(\rm{{EQ}_{tt}-\rm{EQ}_{rr}}\right)+\frac{1}{2}\rm{EQ}_{\theta\theta}=0,$ (18)
and replacing $b$ with $\sqrt{2/\kappa}$, we have
$e^{2\beta(r)}-(1+r\alpha^{\prime}(r))^{2}=0.$ (19)
Substitute Eq. (19) into Eq. (14d) and Eq. (15), surprisingly, we can find that all of the equations become the same equation as
$(\ell-2)r\alpha^{\prime\prime}(r)$ $+\ell r^{2}\alpha^{\prime}(r)^{3}+2(\ell-1)r\alpha^{\prime}(r)^{2}$ (20)
$+2(\ell-2)\alpha^{\prime}(r)=0.$
So in the case of $b=\sqrt{2/\kappa}$, static spherical vacuum solutions can exist, if we could find a solution of Eq. (20).
IV.1 Solutions with constant $\beta$
In this case, from Eq. (19), $\alpha$ can be expressed as
$\alpha(r)=A\ln(\frac{r}{R_{0}}),$ (21)
where $R_{0}$ is a constant. Substituting it into Eq. (20), we can obtain three solutions as $A=0$, $A=-1$, and $A=\frac{2}{\ell}-1$. The case $A=0$ is the flat Minkowski space-time. The case $A=-1$ is ruled out because $g_{rr}=(1+r\alpha^{\prime}(r))^{2}=0$. When $A=\frac{2}{\ell}-1$, we can get
$g_{tt}$ $=-e^{2\alpha}=-\left(\frac{r}{R_{0}}\right)^{2(2/\ell-1)},$ (22)
$g_{rr}$ $=(1+r\alpha^{\prime}(r))^{2}=\frac{4}{\ell^{2}},$
$b_{t}$ $=\sqrt{\frac{2}{\kappa}}e^{\alpha}=\sqrt{\frac{2}{\kappa}}\left(\frac{r}{R_{0}}\right)^{2/\ell-1}.$
Thus the metric of the solution is
$g_{\mu\nu}=\mathrm{diag}\left(-\left(\frac{r}{R_{0}}\right)^{\frac{2(2-\ell)}{\ell}},\frac{4}{\ell^{2}},r^{2},r^{2}\sin^{2}\theta\right).$ (23)
When $\ell=2$, i.e., $\xi=\kappa$, we can see that the above solution degenerates into the flat Minkowski spacetime. And of course, this solution does not admit the limit $\ell\rightarrow 0$. The Kretschmann scalar of this solution is
$K=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{(\ell-2)^{2}(7\ell^{2}-4\ell+8)}{4r^{4}},$ (24)
which means that $r=0$ is a singularity of this solution  [^3] Indeed the other numerator factor $07\ell^{2}-4\ell+8=$ 0 has no real root, therefore the singularity cannot be cured by carefully choosing the value of $\ell$ besides $\ell=2$.. So this solution admits a naked singularity located at $r=0$. The Einstein tensor for the solution Eq. (23) is
$G^{\mu}_{\nu}=\mathrm{diag}\left((\ell^{2}-4)A,-(\ell-2)^{2}A,(\ell-2)^{2}A,(\ell-2)^{2}A\right),$ (25)
where $A=1/4r^{2}$. So we can easily see that when $0\leq\ell\leq 2$, the metric satisfies the dominant energy condition (DEC), and consequently the weak, null, and null dominant energy condition (WEC, NEC, NDEC). Furthermore, the strong energy condition (SEC) is also satisfied in this case.
IV.2 Solutions with non-constant $\beta$
It seems that from Eq. (19) we have two solutions as $1+r\alpha^{\prime}(r)=\pm e^{\beta(r)}$. In fact, this appears because of our choice of coordinates, and both of them provide the same solution. For simplicity, we only consider the case $1+r\alpha^{\prime}(r)=-e^{\beta(r)}$. It is hard to solve Eq. (20) directly. Using
$\alpha^{\prime}(r)=-\frac{1+e^{\beta}(r)}{r}$ (26)
and substitute it into Eq. (20), we get the equation for $\beta(r)$ as
$(\ell-2)r\beta^{\prime}(r)+(e^{\beta}+1)(\ell e^{\beta}+2)=0.$ (27)
Let $u(r)=e^{-\beta(r)}$, we have
$\frac{1}{r}\frac{dr}{du}=\frac{(\ell-2)u}{(u+1)(2u+\ell)},$ (28)
and the solution is
$\frac{r}{R_{0}}=\frac{(2u+\ell)^{\ell/2}}{u+1},$ (29)
where $R_{0}$ is a constant.
Firstly, we consider the case of $\ell\rightarrow 0$ (i.e., $\xi\rightarrow 0$). Then Eq. (29) gives $u=R_{0}/r-1$, leading to
$g_{rr}=e^{2\beta}=\frac{1}{u^{2}}=\left(1-\frac{R_{0}}{r}\right)^{-2}.$ (30)
The solution for $\alpha$ is
$\alpha=\ln\left(1-\frac{R_{0}}{r}\right)+C,$ (31)
thus
$g_{tt}=-e^{2\alpha}=-\left(1-\frac{R_{0}}{r}\right)^{2}$ (32)
up to an overall constant that can be absorbed into a redefinition of $t$. The solution for $B_{\mu}$ is
$b_{t}=\sqrt{\frac{2}{\kappa}}(1-\frac{R_{0}}{r}),$ (33)
and the metric is
$g_{\mu\nu}=\mathrm{diag}\left(-\left(1-\frac{R_{0}}{r}\right)^{2},\left(1-\frac{R_{0}}{r}\right)^{-2},r^{2},r^{2}\sin^{2}\theta\right).$ (34)
Actually, this solution coincides with the extremal Reissner–Nordström black hole solution, while the solution for $B_{\mu}$ is not the same: in the Reissner-Nordström solution, $B_{\mu}=(Q/r,0,0,0)$ which obeys $B_{\mu}(r)=0\ \text{for }r\to\infty$; in our case, however, the temporal component of $B_{\mu}$ at the infinity is a non-zero constant $\sqrt{2/\kappa}$. Nevertheless, the difference is merely a consequence of choosing different boundary conditions for $B_{\mu}$.
For the general cases, from Eq. (29) and Eq. (26), we have
$\frac{d\alpha}{du}=-\frac{1+1/u}{r}\frac{dr}{du}=-\frac{\ell-2}{2u+\ell},$ (35)
and we get
$g_{tt}=-e^{2\alpha}=-C^{2}(2u+\ell)^{-(\ell-2)},$ (36)
where $C$ is a constant. In the coordinate $(t,u,\theta,\phi)$, the $uu-$component of the metric is
$g_{uu}=e^{2\beta}r^{\prime}(u)^{2}=(\ell-2)^{2}R_{0}^{2}\frac{(2u+\ell)^{\ell-2}}{(1+u)^{4}}.$ (37)
Now we try to find a new coordinate system $(t,\rho,\theta,\phi)$, satisfying (1) $g_{tt}\cdot g_{\rho\rho}=-1$, (2) $s\rightarrow r$ when $r\rightarrow\infty$. Condition (1) implies that
$\frac{(\ell-2)^{2}C^{2}R_{0}^{2}}{(1+u)^{4}}u^{\prime}(\rho)^{2}=1,$ (38)
which gives
$u=-1+\frac{(\ell-2)CR_{0}}{\rho}.$ (39)
Combining condition (2) and Eq. (29), we have
$C=(\ell-2)^{\ell/2-1}.$ (40)
We finally arrive at the result that, in the coordinate system $(t,\rho,\theta,\phi)$, the metric can be expressed as
$g_{\mu\nu}=\mathrm{diag}\left(-A(\rho),A(\rho)^{-1},R(\rho)^{2},R(\rho)^{2}\sin^{2}\theta\right),$ (41)
where
$A(\rho)$ $=\left(1-\frac{R_{s}}{\rho}\right)^{2-\ell},$ (42)
$R(\rho)$ $=\left(1-\frac{R_{s}}{\rho}\right)^{\ell/2}\rho,$
and $R_{s}$ is a redefinition of a combination of $\ell$ and $R_{0}$. For $B_{\mu}$, the result is
$b_{t}(r)=\sqrt{\frac{2}{\kappa}}\left(1-\frac{R_{s}}{\rho}\right)^{1-\ell/2}.$ (43)
The ADM mass can be read off from the metric as
$M_{\rm ADM}=\frac{(2-\ell)R_{s}}{2G}.$ (44)
The Kretschmann scalar of this solution is
$K$ $=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$
$=\frac{(\ell-2)^{2}R_{s}^{2}}{4(\rho-R_{s})^{2\ell}\rho^{8-2\ell}}$
×$\phantom{=}\times\left(48\rho^{2}+32(\ell-3)R_{s}\rho+(56-36\ell+7\ell^{2})R_{s}^{2}\right),$ (45)
so when $\ell>0$ and $\ell\neq 2$, $\rho=R_{s}$ is its singularity, and when $\ell<4$ and $\ell\neq 2$, $\rho=0$ is its singularity. When $\ell=2$, this solution degenerates into the flat space-time, just the same as the case discussed in Sec. IV.1. When $\ell\rightarrow 0$, we return to the extremal Reissner–Nordström black hole solution discussed above.
The Einstein tensor for the solution Eq. (41) is
$G^{\mu}_{\nu}=\mathrm{diag}\left((\ell^{2}-4)A,-(\ell-2)^{2}A,(\ell-2)^{2}A,(\ell-2)^{2}A\right),$ (46)
where
$A=\frac{R_{s}^{2}}{4\rho^{4}}\left(1-\frac{R_{s}}{\rho}\right)^{-\ell}.$ (47)
Still, when $0\leq\ell\leq 2$, the metric satisfies the DEC, WEC, NEC, NDEC, and SEC. It is noteworthy that this is also the sufficient condition for the ADM mass Eq. (44) to be non-negative.
From the solutions Eq. (23) and Eq. (41), we know that there is no Schwarzschild-like solution when the bumblebee field obtains a time-like VEV as $b=\sqrt{2/\kappa}$.
V Discussion and Summary
V.1 The stability of the solutions with VEV $b=\sqrt{2/\kappa}$
Here we argue that the solutions with the VEV $b=\sqrt{2/\kappa}$ discussed in Sec. IV are generally unstable, and thus complete one of our main results: there is no non-trivial static spherical solution for bumblebee gravity with time-like VEVs.
First, there is a necessary condition for the stability of the bumblebee model: the Hamiltonian is bounded from below [11], and thus the functional form of the potential is constrained. To our knowledge, stable models are only known for $V\propto X\equiv B_{\mu}B^{\mu}+sb^{2}$ and $V\propto M(n,2,X/\mu^{2})-1$ ($n\geq 3$ for time-like VEVs), where $\mu^{2}$ is an energy scale of the theory and $M(\alpha,\beta,z)$ is the confluent hypergeometric (Kummer’s) function [12, 13]. However both types of potentials only statisfy the condition $V(0)$ while violate $V^{\prime}(0)=0$, for which one cannot solve the equations of motions by assuming a fixed vacuum expectation value of $B_{\mu}$. On the other hand, the conditions in Eq. (3) can most easily be realized by polynomial functions $V\propto X^{n},\ n>1$, and it is shown that in general there is no lower bound on the Hamiltonian [11, 14, 15] for such kind of potentials. This observation then implies the instability of the solutions with $b=\sqrt{2/\kappa}$.
Second, the VEV condition $b=\sqrt{2/\kappa}$ is such a strict requirement, that any deviation, regardless of how small it is, will ultimately cause the non-trivial solutions degenerate into the Minkowski space-time. This is evident if we observe that Eq. (17) forces $\alpha^{\prime}(r)$ to be zero once $\kappa b^{2}\neq 2$ except at $r=0$. Although the seemingly numerical coincidence is maintained once we set $b=\sqrt{2/\kappa}$ at the classical level, quantum corrections would unavoidably shift the VEV of the bumblebee field $B_{\mu}$, and lead the space-time solutions to decay into the flat space-time. If we focus on the bumblebee field sector, similar to the situation in quantum field theories, loop corrections in general could modify the form of the potential $V\to V_{eff}$, for which the location of the minimum can be different from 00 as in Eq. (3). An immediate consequence is that the VEV gets modified correspondingly $b\to b_{eff}\neq\sqrt{2/\kappa}$. Put differently, the non-trivial solutions with $b=\sqrt{2/\kappa}$ are not stable once quantum fluctuations are taken into account. Of course, this argument fails if, because of some unknown mechanisms of the theory of QG, $b$ is fine-tuned to be exactly $\sqrt{2/\kappa}$, which holds to any order of quantum corrections; or, the gravity coupling constant $\kappa$ would also change by QG corrections such that $b_{eff}=\sqrt{2/\kappa_{eff}}$ throughout. However, even without full understanding of the QG theories, we believe that these mechanisms preventing the non-trivial solutions from decay are quite impossible.
In summary, even each of the two arguments above may be invalid once a complete theory of quantum gravity is taken into account, they, put together, still strongly imply the non-existence of a non-trivial solution with any value of $b$ in the case of time-like VEVs studied in our work.
V.2 Violation of the weak cosmic censorship conjecture
The weak cosmic censorship conjecture (WCCC), proposed by Sir Roger Penrose [16], serves as a method to save the predictability of general relativity [17]. However, our solutions in this work provide explicit examples that counter this conjecture. In fact, the solution Eq. (23) and the corresponding Kretschmann scalar Eq. (24) indicate that there is a singularity located at $r=0$, but no horizon exists in this case, meaning the existence of a naked singularity, and the violation of the WCCC consequently. Nonetheless, we have argued that the solutions in Sec. IV are unstable, therefore the violation of the WCCC appears to be harmless, since the solutions would finally decay into the flat space-time.
V.3 Summary
In this work, we study the static spherical vacuum solution for bumblebee gravity with time-like VEVs as $b_{\mu}=(b_{t}(r),0,0,0)$. The results suggest that for general VEVs, there is no consistent non-trivial solution if $b\neq\sqrt{2/\kappa}$. With $b=\sqrt{2/\kappa}$, we find two non-trivial solutions as Eq. (23) and Eq. (41), none of which is Schwarzschild-like. And we argue that these solutions are unstable, which comes to the rescue when the predictability of the theory is ruined because of the existence of naked singularities. Since there are no Schwarzschild-like solutions for bumblebee gravity with time-like VEVs, and only the flat space-time solution is stable, we probably do not need to consider the case of time-like VEVs when related to real physics in future researches. However, the newly obtained solutions give rise to further questions, such as what is the reason for the existence of a solution when $b=\sqrt{2/\kappa}\simeq E_{Pl}$.
Acknowledgements.
The authors are grateful for valuable discussions with Hanlin Song and Peixiang Ji. This work was supported in part by the National Science Foundation of China under Grant No. 12347101. HL was also supported by the start-up fund of Chongqing University under No. 0233005203009, and JZ was supported by the start-up fund of Chongqing University under No. 0233005203006.